OpenStudy (anonymous):

Find a cofunction with the same value as the given expression. csc 52° A. sec 52° B. sec 38° C. sin 52° D. sec 142°

OpenStudy (anonymous):

B

OpenStudy (anonymous):

csc (x) = sec ( 90 - x) so csc(52) = sec (90 - 52) = sec (38)